Entries for Sale of Fixed Asset

Equipment acquired on January 5 at a cost of $117,400, has an estimated useful life of 15 years, has an estimated residual value of $7,450, and is depreciated by the straight-line method.

a. What was the book value of the equipment at December 31 the end of the fourth year?
$

b. Assuming that the equipment was sold on April 1 of the fifth year for 79,642.

1. Journalize the entry to record depreciation for the three months until the sale date. Round your answers to the nerest whole dollar if required.

2. Journalize the entry to record the sale of the equipment. If an amount box does not require an entry, leave it blank. Do not round intermediate calculations.
Answer #1
1
Cost 117400
Less: Residual value 7450
Depreciable cost 109950
Divide by useful life 15
Annual depreciation 7330
Depreciation for 4 years 29320 =7330*4
Cost 117400
Less: Depreciation for 4 years 29320
Book value of the equipment 88080
b
Depreciation Expense-Equipment 1833 =7330/12*3
         Accumulated Depreciation-Equipment 1833
2
Cash 79642
Accumulated Depreciation-Equipment 31153 =29320+1833
Loss on Sale of Equipment 6605
         Equipment 117400
Note: Depreciation for 3 months is $1832.5. It has been rounded off to 1833
Answer might vary +1 due to rounding off